# Enabling Low-Cost Secure Computing on Untrusted In-Memory Architectures

### v1

@@ 1 Introduction
    Empirical Evaluation on Real Hardware: To the best of our knowledge, this is the first work to evaluate MPC-based secure computation on real-world PIM hardware (UPMEM). Our implementation and evaluation are publicly available [^1] https://zenodo.org/records/14736864, allowing other researchers to build upon our findings.
@@ 3.3 Trusted Execution Environment Extension
Trusted Execution Environment (TEE) is a secure area inside a main processor. It guarantees code and data loaded inside are protected with respect to confidentiality and integrity. In the scope of TEE, memory encryption technologies play a vital role in ensuring data security. AES is a commonly used memory encryption technology [38, 10]. By using AES, data that leaves the processor to be stored in the memory is encrypted, thus protecting the confidentiality of the information against potential attackers. For the integrity of data - ensuring that it has not been altered or tampered with, a data structure known as a Merkle Tree is often employed [38, 10]. By hashing the data and constructing a tree of hashes, the integrity of large sets of data can be verified efficiently and securely.
@@ 6.1 Workload
Implementation of MLP: MLP [31] is a class of neural networks with at least three layers. Typically, each layer consists of a GEMV kernel followed by the activation function (ReLU) [25]. In our baseline UPMEM implementation of MLP inference [25], each layer of the MLP involves a weight matrix and an input vector. We consider the vector inputs to be private and weights to be public.
@@
In the implementation of logistic regression training, we adhere to the approach outlined in our UPMEM-Insecure baseline study [24]. Initially, the input is distributed row-wise across PIM cores and tasklets. The CPU performs the final reduction step in each iteration and updates the model parameters.
@@
Implementation of Linear Regression: Linear regression [64], is a statistical technique used to model the optimal relationship between a dependent variable and one or more independent variables using a linear equation. Like logistic regression, the training involves multiple iterations to find optimal parameters. Each iteration is identical to logistic regression except for the activation function. In the baseline implementation of linear regression [24], the input is distributed across PIM cores and tasklets. Partial results are then merged within DPUs and CPU before updating the weights. Since we execute both dot product and gradient descent on UPMEM, some additional communication between CPU and PIM, after dot products and before gradient descent, is necessary.
@@ Acknowledgments
This work is funded, in part, by NSF Career Award #2339317, the initial complement fund from UCR, and the Hellman Fellowship from the University of California. We appreciate the use of hardware resources provided by the UPMEM. We thank Prof. Nael Abu-Ghazaleh for his insightful comments.

### v2

@@ 1 Introduction
    Empirical Evaluation on Real Hardware: To the best of our knowledge, this is the first work to evaluate MPC-based secure computation on real-world PIM hardware (UPMEM). Our implementation and evaluation are publicly available [^1] https://zenodo.org/records/14736863 , allowing other researchers to build upon our findings.
@@ 3.3 Trusted Execution Environment Extension
Trusted Execution Environment (TEE) is a secure area inside a main processor. It guarantees code and data loaded inside are protected with respect to confidentiality and integrity. In the scope of TEE, memory encryption technologies play a vital role in ensuring data security. AES is a commonly used memory encryption technology [41, 10, 43]. By using AES, data that leaves the processor to be stored in the memory is encrypted, thus protecting the confidentiality of the information against potential attackers. For the integrity of data - ensuring that it has not been altered or tampered with, a data structure known as a Merkle Tree is often employed [41, 10]. By hashing the data and constructing a tree of hashes, the integrity of large sets of data can be verified efficiently and securely.
@@ 6.1 Workload
Implementation of MLP: MLP [34] is a class of neural networks with at least three layers. Typically, each layer consists of a GEMV kernel followed by the activation function (ReLU) [27]. In our baseline UPMEM implementation of MLP inference [27, 28, 24], each layer of the MLP involves a weight matrix and an input vector. We consider the vector inputs to be private and weights to be public.
@@
In the implementation of logistic regression training, we adhere to the approach outlined in our UPMEM-Insecure baseline study [26, 23]. Initially, the input is distributed row-wise across PIM cores and tasklets. The CPU performs the final reduction step in each iteration and updates the model parameters.
@@
Implementation of Linear Regression: Linear regression [68], is a statistical technique used to model the optimal relationship between a dependent variable and one or more independent variables using a linear equation. Like logistic regression, the training involves multiple iterations to find optimal parameters. Each iteration is identical to logistic regression except for the activation function. In the baseline implementation of linear regression [26, 23], the input is distributed across PIM cores and tasklets. Partial results are then merged within DPUs and CPU before updating the weights. Since we execute both dot product and gradient descent on UPMEM, some additional communication between CPU and PIM, after dot products and before gradient descent, is necessary.
@@ Acknowledgments
This work is funded, in part, by NSF Career Award #2339317, NSF #2235398, the initial complement fund from UCR, and the Hellman Fellowship from the University of California. We appreciate the use of hardware resources provided by the UPMEM. We thank Prof. Nael Abu-Ghazaleh for his insightful comments.
@@
Appendix A Artifact Appendix
A.1 Abstract
Our artifact contains all necessary source codes and scripts for evaluating our proposed security scheme. This paper leverages Multi-Party Computation (MPC) techniques, specifically arithmetic secret sharing, and Yao’s garbled circuits, to securely outsource bandwidth-intensive computation to PIM. Additionally, we employ precomputation optimizations to prevent the CPU’s portion of the MPC from becoming a bottleneck. We provided all the source codes and scripts for evaluating our scheme using UPMEM, the first publicly available PIM, over four data-intensive applications: Multilayer Perceptron inference (MLP), Deep Learning Recommendation Model inference (DLRM), linear regression training, and logistic regression training. This artifact allows researchers to reproduce our results, explore this area further, and expand our work. With this artifact, researchers can regenerate the performance results for UPMEM-Precompute-C(V), UPMEM-Runtime-C(V), and UPMEM-Runtime-A(2Y)-C(V), as presented in figures 13 to 18.
A.2 Description & Requirements
A.2.1 Security, privacy, and ethical concerns
Our research does not compromise user privacy or safety, and all data used in our experiments is either synthetic or randomly generated. Therefore, no Personally Identifiable Information (PII), sensitive data, or human subjects were involved.
A.2.2 How to access
The code is open-source and hosted on a GitHub public repository to facilitate reuse. Additionally, it is archived on Zenodo. We encourage other researchers to explore, reproduce, and modify our code, provided they give appropriate credit. We always recommend using the latest Zenodo release via the provided concept DOI.
A.2.3 Hardware dependencies
To evaluate our method, we utilize UPMEM PIM hardware, which primarily consists of standard DDR4-2400 DIMMs integrated with DPUs. Our setup includes 20 PIM-enabled DIMMs, providing a total of 160 GB of MRAM and 2560 DPUs working in parallel at a clock frequency of 350 MHz. The host server for the UPMEM system is equipped with a 2-socket Intel Xeon Silver 4110 CPU. To accurately reproduce our results, access to the actual hardware is necessary. UPMEM’s PIM data centers are accessible upon request at https://www.upmem.com/developer.
A.2.4 Software dependencies
All the implemented applications require the UPMEM SDK, which can be installed based on the hardware specifications and is accessible at https://sdk.upmem.com. For dependencies related to MLP, DLRM, logistic regression, and linear regression, please refer to our baseline implementations as our work is built upon them.
A.2.5 Benchmarks
Our scheme is evaluated using MLP inference, DLRM inference, logistic regression training, and linear regression training. To evaluate our implementation, we use randomly generated inputs.
A.3 Set-up
A.3.1 Installation
To run this artifact on your local device, the UPMEM SDK must first be installed, which is available at https://sdk.upmem.com. However, as previously mentioned, reproducing our results requires access to real hardware (20 UPMEM PIMs) rather than the simulator. Once the UPMEM SDK is installed, the artifact can be downloaded from GitHub and Zenodo.
A.3.2 Basic Test
Script run_functionality.sh, in the root directory, can be used to perform a simple functionality check. This file executes a basic test on all the applications sequentially and outputs execution time. To run a specific application, the run_functionality.sh script in the corresponding folder can be executed.
A.4 Evaluation workflow
A.4.1 Major Claims
Our major claim is as follows:
(C1):
Compared to a secure CPU implementation, our framework achieves speedups of 14.66×14.$6$, 9.80×9.80$\times$9.80 ×, 2.64×2.64$\times$2.64 ×, and 5.85×5.85$\times$5.85 × for MLP inference, DLRM inference, Linear Regression training, and Logistic Regression training, respectively. This is proven by the experiment (E1) described in Section 7.1 whose results are illustrated in Figures 14, 15, 16, 18.
A.4.2 Experiments
(E1):
[30 human-minutes + 2.5 compute-hour + 32GB disk]:
How to:
Our results can be reproduced by following the three steps below.
Preparation:
Install the UPMEM SDK, as described in Section A.2.4, then clone our artifact.
Execution:
Script ./run_reproduce.sh, in the root directory, can be used to regenerate our results. This file executes all the applications with our configuration sequentially and outputs execution time for UPMEM-Precompute-C(V), UPMEM-Runtime-C(V), and UPMEM-Runtime-A(2Y)-C(V), as presented in Figures 13 to 18. To reproduce the results for a specific application, the ./run_reproduce.sh script in the corresponding folder can be executed.
This script compiles and links the host and DPU source codes for a specific number of DPUs and tasklets. There is a Makefile for each application that facilitates the compilation and linking of the source codes.
Results:
After running our script, the final execution time is reported. Table 2 explains the different notations used for manually determining execution times. The execution time is calculated using the formula provided below:
$PIM\ Time=CPU-DPU+PIM\ kernel+DPU-CPU$
$Kernel\ time=Max[CPU\ Time,PIM\ Time]$
$Execution\ time=Kernel\ time+Merge+Verification$
Table 2: List of timing notations
Timing notations Explanation
$CPU-DPU$ CPU to DPUs transfer time
$DPU-CPU$ DPUs to CPU transfer time
$PIM\ kernel$ Kernel execution time on DPUs
$CPU\ Time$ Kernel execution time on CPU
$Kernel\ time$ The maximum of CPU Time and PIM Time
$Merge$ Merging time of CPU and PIM results
$Verification$ Verification time
A.5 Version
Based on the LaTeX template for Artifact Evaluation V20231005. Submission, reviewing and badging methodology followed for the evaluation of this artifact can be found at https://secartifacts.github.io/usenixsec2025/.